OpenStudy (rainstark):

Need help with this Lim (sin2x+tan3x)/x when x->0+ Result = 5

OpenStudy (sooobored):

sen? or sin?

OpenStudy (rainstark):

sin, sorry

OpenStudy (sooobored):

\[\frac{\sin 2x+\tan 3x}{x}\]

OpenStudy (rainstark):

yeah

OpenStudy (sooobored):

\[\frac{\sin 2x}{x} + \frac{\tan 3x}{x}\]

OpenStudy (sooobored):

\[\frac{2\sin 2x}{2x} + \frac{3 \tan 3x}{3x}\]

OpenStudy (sooobored):

\[\lim_{x \rightarrow 0} \frac{sinx}{x} =1\]

OpenStudy (sooobored):

\[\tan x = \frac{\sin x}{\cos x}\]

OpenStudy (sooobored):

\[2\frac{\sin 2x }{x} + \frac{3}{\cos{3x}} \frac{\sin 3x}{3x}\]

OpenStudy (sooobored):

crap i made a typoooooo

OpenStudy (sooobored):

pretend its divided by 2x instead of x

OpenStudy (sooobored):

\[\frac{\sin 2x+\tan 3x}{x} = 2 \frac{\sin 2x}{2x} + \frac{3}{\cos 3x} \frac{\sin 3x}{3x}\]

OpenStudy (sooobored):

so, can you do the limit now? \[\lim_{x \rightarrow 0^+} 2 \frac{\sin 2x}{2x} + \frac{3}{\cos 3x}\frac{\sin 3x}{3x}\]

OpenStudy (rainstark):

3/ 3cos(0) = 3/3 = 1 ?

OpenStudy (rainstark):

2+3 = 5 ?

OpenStudy (sooobored):

the lim of sin 2x /2x is 1 the lim of sin 3x/ 3x is 1 the lim of 3/ cos 3x = 3/1 = 3

OpenStudy (sooobored):

its 3/ cos 3x not 3/3cos3x

OpenStudy (sooobored):

and then yes, 2+3 = 5

OpenStudy (rainstark):

OHHH, sorry. I am a lil bit sleepy. Thank you so much. You saved my life

OpenStudy (sooobored):

when in doubt, see if any trig identities will make the problem easier

OpenStudy (rainstark):

Thank you so muuuuch. I hope you are having a great night <3